# Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

### v2

@@ Abstract
Uncovering hidden symbolic laws from time series data, as an aspiration dating back to Kepler’s discovery of planetary motion, remains a core challenge in scientific discovery and artificial intelligence. While Large Language Models show promise in structured reasoning tasks, their ability to infer interpretable, context-aligned symbolic structures from time series data is still underexplored. To systematically evaluate this capability, we introduce SymbolBench, a comprehensive benchmark designed to assess symbolic reasoning over real-world time series across three tasks: multivariate symbolic regression, Boolean network inference, and causal discovery. Unlike prior efforts limited to simple algebraic equations, SymbolBench spans a diverse set of symbolic forms with varying complexity. We further propose a unified framework that integrates LLMs with genetic programming to form a closed-loop symbolic reasoning system, where LLMs act both as predictors and evaluators. Our empirical results reveal key strengths and limitations of current models, highlighting the importance of combining domain knowledge, context alignment, and reasoning structure to improve LLMs in automated scientific discovery.
@@
Zewen Liu1, Juntong Ni1, Xianfeng Tang2, Max S.Y. Lau1, Wenpeng Yin3, Wei Jin1 1Emory University, 2Amazon, 3Penn State University {zewen.liu, juntong.ni, msy.lau, wei.jin}@emory.edu, xianfengtang@outlook.com, wenpeng@psu.edu
@@ 1 Introduction
To realize these objectives, we present three primary contributions.
(I) SymbolBench, a comprehensive benchmark designed to rigorously evaluate the symbolic reasoning capabilities of LLMs over time series with rich real-world contextual descriptions. It uniquely spans three core tasks covering major types of time series data: (a) Multivariate symbolic regression for continuous data to recover complex equations like coupled Ordinary Differential Equations (ODEs). (b) Boolean network inference Zhang et al. (2024) for discrete systems to identify logical rules. (c) Causal discovery Assaad et al. (2022) for multivariate data to uncover structured causal graphs. Each task includes challenging, real-world examples from domains such as biology, physics, and healthcare, with varying dimensionality and difficulty. To ensure both quality and coverage, we curate representative subsets from large databases, yielding a benchmark that is broader, more challenging, and more balanced than prior efforts.
(II) Unified Symbolic Reasoning Framework that enables the context-aware and knowledge-rich LLMs to play the dual role of predictors and judges in the process of hypothesis generation, testing, and refinement, while optionally including efficient genetic programming tools in a hybrid way.
(III) Critical empirical insights into the current strengths and limitations of LLMs and MLLMs in temporal symbolic reasoning. (a) LLMs surpass traditional baselines on multivariate symbolic regression and causal discovery, but fall short on Boolean network inference; (b) LLMs are able to perform a certain level of reasoning on the three tasks, with properly increasing test-time compute bringing moderate improvement. (c) Introducing context not only improves performance but also potentially guides the selection of generalizable symbolic structures. (d) Combining LLMs in complementary roles with genetic programming further boosts performance.
@@ 2 Related Work
LLM Symbolic Reasoning. Due to the strong in-context learning and reasoning ability of LLMs that allow them to adapt to various tasks, they are able to perform logical inferenceAhn et al. (2024); Wang and Chen (2023), and temporal symbolic reasoning Fang et al. (2024). Despite the strong ability of LLMs, current research for SR tasks only applies them as function generators without a reasoning process. Though the generated functions may achieve a high fitting score, the reasoning process remains unknown to us and may fail to align with the context, with little real-world meaning. In addition, the current research focus has also skewed toward algebraic equations, with logical rules or causal relations remaining underexplored.
@@ 3 SymbolBench Dataset
To rigorously evaluate the reasoning abilities of LLMs on time series related science discovery, we introduce SymbolBench, a curated benchmark that aims to uncover symbolic structures from time series. The dataset spans diverse domains (e.g., physics, biology) and is structured around three core categories of symbolic structures. More details of the dataset are provided in Appendix B.
Coupled Differential Equations (CDEs). CDEs represent dynamic systems, yielding continuous and multivariate time series data. Let $x_{i}$ denote the $i$-th variable of a multivariate time series, a coupled differential equation can be described as the following: $\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{n}),i=1,...,n$, where the symbolic structures $f_{i}$ describe dynamic interactions among state variables. The corresponding time series data consists of numerical solutions $\{x_{i}(t)\}$ over time, generated from initial conditions and system parameters. This setting reflects real-world dynamical systems in physics and engineering, where the complexity arises from variable interdependence. While the previous benchmark dataset, ODEbench d’Ascoli et al. (2023), contains coupled ODEs from 1 dimension to 4 dimensions, the number of samples is small, and the class of dimensions is heavily imbalanced, with only three 4-dimensional ODEs. In this study, we further enrich ODEbench with more high-dimensional ODEs and provide a balanced dataset with over 156 samples. Each sample is accompanied by the variable descriptions and the domain name if available, as shown in Appendix B.
@@
Figure 1: Iterative refinement framework. Candidate proposals are generated using either an LLM-as-Predictor or genetic programming operations. Each round of candidates undergoes quantitative and qualitative evaluation via validation tools and an LLM-as-Judge. Scored candidates are stored in a history pool, and a context manager decides contextual information for the next round.
@@ 4.1 Proposal Generation
LLM-as-Predictor. To leverage the knowledge embedded in LLMs, recent studies have explored the direct generation of equations. In this work, we frame this approach as LLM-as-Predictor. To thoroughly assess both the capabilities of LLMs and the impact of iterative refinement, external context, and chain-of-thought reasoning, we introduce four distinct prompting strategies to evaluate the performance of LLMs/MLLMs comprehensively. (a) Naive prompt instructs the model to generate mathematical expressions without providing any contextual or historical information. (b) Base prompt builds upon the Naive prompt by incorporating previously generated expressions from the history pool, filtered by the context manager. (b) Context prompt further enhances the Base prompt by adding relevant contextual information, such as variable descriptions. (c) CoT prompt extends the Context prompt by enabling the model to perform step-by-step reasoning. The reasoning process is also recorded and subjected to qualitative verification. Details of all prompts are provided in Appendix I.
Hybrid Method. While the generation of symbolic structures is central to the pipeline, this part can also be addressed using operations from traditional genetic programming. In the context of CDEs and BNs, this involves applying operations like mutation and crossover on expression trees to produce new candidates. Since LLMs may still contribute to other components of the framework, e.g., verification, we phrase such a combination as the Hybrid Method, and further discuss it in Appendix F.
@@ 4.2 Verification
Verification in symbolic structure discovery assesses three key aspects: numerical fitness, symbolic fitness, and context alignment. While numerical fitness can be quantified, symbolic fitness and context alignment are harder to measure without ground truth. To address this, we introduce a dual-verification strategy combining standard quantitative metrics with rubric-based qualitative evaluation by LLMs-as-Judges.
@@ 4.2.1 Quantitative Verification
(a) For CDEs, we simulate time series data using the predicted functions and assess their numerical fitness using the averaged coefficient of determination: $R^{2}=1/D\sum_{j}^{D}(1-\frac{\sum_{t=1}^{T}\|x^{t}_{j}-\hat{x}^{t}_{j}\|^{2}}{\sum_{t=1}^{T}\|x^{t}_{j}-\bar{\mathbf{x_{j}}}\|^{2}}),$ where $x^{t}_{j}$ denotes the ground truth at time $t$ for time series $j$, $\hat{x}^{t}_{j}$ the model prediction, and $\bar{\mathbf{x}_{j}}$ the empirical mean of the observed data. We also evaluate expression complexity based on symbolic structure, defined as the number of operations. (b) For Boolean networks, we simulate transitions from the predicted Boolean networks and evaluate predictive performance using the macro-averaged F1 score over $T$ transitions: $\mathrm{F1}=\frac{1}{D*T}\sum_{j=1}^{T}\sum_{i=1}^{D}\frac{2\cdot\mathrm{TP}^{j}_{i}}{2\cdot\mathrm{TP}^{j}_{i}+\mathrm{FP}^{j}_{i}+\mathrm{FN}^{j}_{i}},$ where $\mathrm{TP}^{j}_{i}$, $\mathrm{FP}^{j}_{i}$, and $\mathrm{FN}^{j}_{i}$ are the true positives, false positives, and false negatives for node $i$ at transition $j$, respectively. Expression complexity is computed analogously to CDEs. (c) For structural causal models, where the true data-generating process is typically unknown, we adopt a Conditional Independence score (CI-score) as a proxy for structural fidelity. This metric captures the strength of dependence between each child node and its parents, conditioned on the remaining parents. Let the parent set of $X^{i}$ defined as $P_{i}=\{(\mathbf{x}_{j},\ell)|j\in D,\ell\in\ell_{\max}\}$ where $\ell_{\max}$ is the maximum time lag. For each directed edge $(\mathbf{x}_{j},\ell)\to i$ in the edge set $E$, we compute the partial correlation: $\scriptstyle r_{j\to i}=\operatorname{Corr}\!\left(X^{t-\ell}_{j},\;X^{t}_{i}\;\middle|\;\left\{x^{t-\ell^{\prime}}_{k}:(\mathbf{x}_{k},\ell^{\prime})\in P_{i},\;k\neq j\right\}\right).$ The CI-score for a candidate graph $\mathcal{G}$ is then given by: $\mathrm{CI\text{-}score}(\mathcal{G})=\frac{1}{|E|}\sum_{(\mathbf{x}_{j},\ell)\to i\in E}\left|r_{j\to i}\right|.$ A lower CI-score suggests stronger conditional independence and thus better structural plausibility.
@@ 4.2.2 Qualitative Verification
To assess the symbolic quality of the generated expressions, we employ a rubric-based scoring system ranging from 1 (poor) to 5 (excellent). Each qualitative score is assigned by an LLM acting as a judge Hao et al. (2024), using a standardized evaluation rubric. The rubric consists of four core criteria: (a) Context alignment: Alignment with provided time series data and contextual descriptions; (b) Scientific plausibility: Alignment with possible physical laws or domain-specific constraints; (c) Conciseness and clarity: The readability and succinctness of the reasoning path; (d) Logical coherence: The consistency and step-by-step soundness of the derivation process.
4.3 Context Manager
The Context Manager is responsible for delivering relevant contextual information to the Proposal Generator during iterative refinement. Based on the chosen refinement strategy, the Context Manager selects and provides appropriate context for the next round of proposal generation. Throughout the process, the Context Manager maintains a history pool $\mathcal{H}$, which stores previously generated expressions, reasoning paths, and their associated verification scores in a structured DataFrame, while removing duplicates. In this study, we adopt a naive ranking-based strategy by only selecting the top-$k$ highest-scoring candidates for further refinement.
@@ 5 Experiment
5.1 Experimental Setup
5.2 Experimental Results
Across all three tasks, we select LLMs with various sizes and architectures, including Qwen2.5-14B Team (2024), Llama-3.2B Dubey et al. (2024), Mathstral-7B Jiang et al. (2023), GPT-4o-mini Achiam et al. (2023), and ChatTS-14B Xie et al. (2024). For GPT-4o-mini and ChatTS-14B, we use inputs with visual and temporal modalities. We present the results in Table 2,  3, and  4.
Obs. 1: LLMs demonstrate superior capability compared to baselines on CDEs and SCMs datasets, while failing to compete against baselines in Boolean network inference. For Boolean network inference, while LLMs consistently achieve positive bookmaker informedness scores, indicating performance better than random guessing, the genetic programming-based model, LogicGep, significantly outperforms all evaluated LLMs across nearly all evaluation metrics in both ID and OOD scenarios. This could be explained by both the symbolic and numerical fitting process. (a) Symbolic fitting Unlike CDEs, BN inference does not involve coefficient optimization. Compared with CDEs, even if the symbolic structure is not precisely correct, adjusting coefficients can still yield good numerical accuracy. Compared with SCMs, SCM inference focuses only on discovering causal relationships, which is a less stringent goal than recovering the full dynamics as in BN inference. (b) Numerical fitting: CDEs and SCMs are inferred from continuous time series, which form a consistent chains of state transitions, whereas BN inference uses state transitions with various initial conditions, forming graphs of state transitions, as shown in Appendix B and  H.1, which can be hard to identify trends and summarize patterns.
Obs. 3: Chain-of-thought prompting does not consistently improve performance. Further introducing CoT prompting does not lead to consistent gains, especially on the CDEs dataset. An exception is ChatTS-14B, likely due to its specialized time series reasoning capabilities. While increasing test-time compute has been shown to improve outcomes in various tasks Snell et al. (2024), and CoT can contribute to this effect, similar limitations have been observed in more complex tasks such as those in SciBench Wang et al. (2023).
Obs. 4: Providing problem contexts improves LLM performance. Across all three tasks, compared to the Naive prompt, LLMs demonstrate the ability to leverage the provided context, resulting in higher numerical performance. In this setting, context functions as a form of conditioning, helping to constrain the solution space and guide the model toward more accurate and relevant inferences with a faster convergence rate, as shown in Appendix G.
Table 2: Symbolic regression performance for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Yellow, Orange, and Cyan mark the first, second, and third place, respectively.

### v3

@@ Abstract
Uncovering hidden symbolic laws from time series data, as an aspiration dating back to Kepler’s discovery of planetary motion, remains a core challenge in scientific discovery and artificial intelligence. While Large Language Models show promise in structured reasoning tasks, their ability to infer interpretable, context-aligned symbolic structures from time series data is still underexplored. To systematically evaluate this capability, we introduce SymbolBench [^1] https://anonymous.4open.science/r/SymbolBench-6F40, a comprehensive benchmark designed to assess symbolic reasoning over real-world time series across three tasks: multivariate symbolic regression, Boolean network inference, and causal discovery. Unlike prior efforts limited to simple algebraic equations, SymbolBench spans a diverse set of symbolic forms with varying complexity. We further propose a unified framework that integrates LLMs with genetic programming to form a closed-loop symbolic reasoning system. Our empirical results reveal key strengths and limitations of current models, highlighting the importance of combining domain knowledge, context alignment, and reasoning structure to improve LLMs in automated scientific discovery.
@@
Zewen Liu1, Juntong Ni1, Xianfeng Tang2, Max S.Y. Lau1, Wenpeng Yin3, Wei Jin1 1Emory University, 2Amazon, 3Penn State University {zewen.liu, juntong.ni, msy.lau, wei.jin}@emory.edu, xianfengtang@outlook.com; wenpeng@psu.edu
@@ 1 Introduction
Figure 1: Symbolic structure discovery from time series.
@@
To realize these objectives, we present three primary contributions. (I) SymbolBench, a comprehensive benchmark designed to rigorously evaluate the symbolic reasoning capabilities of LLMs over time series with rich real-world contextual descriptions. It uniquely spans three core tasks covering major types of time series data: (a) Multivariate symbolic regression for continuous data to recover complex equations like coupled Ordinary Differential Equations (ODEs). (b) Boolean network inference Zhang et al. (2024) for discrete systems to identify logical rules. (c) Causal discovery Assaad et al. (2022) for multivariate data to uncover structured causal graphs. Each task includes challenging, real-world examples from domains such as biology, physics, and healthcare, with varying dimensionality and difficulty. To ensure both quality and coverage, we curate representative subsets from large databases, yielding a benchmark that is broader, more challenging, and more balanced than prior efforts. (II) Unified Symbolic Reasoning Framework that enables the context-aware and knowledge-rich LLMs to play the dual role of predictors and judges in the process of hypothesis generation, testing, and refinement, while optionally including efficient genetic programming tools in a hybrid way. (III) Critical empirical insights into the strengths and limitations of LLMs and MLLMs in temporal symbolic reasoning: (a) LLMs outperform traditional baselines on multivariate symbolic regression and causal discovery but lag in Boolean network inference; (b) they demonstrate task-level reasoning, with increased test-time compute yielding moderate gains; (c) contextual information enhances both performance and the discovery of generalizable symbolic structures; (d) integrating LLMs with genetic programming in complementary roles further improves results.
@@ 2 Related Work
LLM Symbolic Reasoning. LLMs demonstrate strong in-context learning and logical reasoning abilities Ahn et al. (2024); Wang and Chen (2023), and have been applied to time series analysis Fang et al. (2024). However, most prior work, such as ChatTS Xie et al. (2024), LLM-ABB Carson et al. (2024), and TemporalGPT Zhang et al. (2025), focuses on the compression or reasoning at the linguistic level, summarizing trends, or producing descriptive insights, rather than uncovering hidden symbolic structures. By contrast, our tasks target recovery of scientific equations, logical rules, and causal graphs, which demand abstraction, contextual alignment, and explanatory power. Existing attempts in symbolic regression often reduce LLMs to function generators Merler et al. (2024); Li et al. (2024); Shojaee et al. (2024), yielding high data fit but low context-alignment, and largely overlook symbolic forms beyond algebraic functions Zhang et al. (2024); Assaad et al. (2022), e.g., Boolean networks Zhang et al. (2024) and causal relations Assaad et al. (2022).
@@ 3 SymbolBench Dataset
To rigorously assess symbolic reasoning over time series, we curate SymbolBench by systematically constructing datasets that combine real-world scientific systems with carefully controlled simulation and contextual annotation. Across all tasks, our construction follows a unified pipeline: (i) identify and extract well-formed ground-truth symbolic structures from established scientific repositories, (ii) generate time series trajectories through simulation under varied initial conditions, (iii) enrich each sample with available domain and variable-level metadata, and (iv) balance dataset distributions to avoid bias toward low-dimensional or trivial cases. Eventually, the goal of each task is to recover the symbolic structure from the corresponding time series. This design allows SymbolBench to address three key dataset objectives: real-world relevance, task difficulty, and scale and balance.
(a) Coupled Differential Equations (CDEs). We extend ODEBench d’Ascoli et al. (2023) by incorporating additional high-dimensional systems, addressing its imbalance of a few four-dimensional examples. Using the Physiome library Yu et al. (2011), which includes over 500 cell-dynamics models, we exclude systems with discrete states, external inputs, or excessively long coefficient lists, focusing on continuous dynamics with symbolic structural complexity. For each selected system, we integrate the governing equations under multiple initial conditions and parameter settings to generate diverse continuous trajectories. Each sample includes textual metadata (variable descriptions and domain labels, e.g., physics, biology)for contextual grounding. The final CDE dataset comprises over 150 balanced samples spanning one- to four-dimensional systems across diverse domains such as physical mechanics, electronics, social dynamics, epidemiology, calcium signaling, neurobiology, and cardiovascular circulation. Each sample uses 150 data points for fitting and evaluation.
(b) Boolean Networks (BNs). To represent discrete dynamical systems, we extract Boolean networks from BioDivine Pastva et al. (2023), a repository of biologically inspired gene regulatory and signaling networks. Each variable is represented as a binary node whose state evolves via logical update rules (See Table 8 in the appendix). We select networks with fewer than 20 variables and filter out those with trivial dynamics (e.g., all states collapsing to a single attractor in one step). We simulate temporal trajectories from given initial conditions, yielding sequences of binary state transitions. Independent variables are randomly assigned values to ensure variability. Alongside the logical rules, we provide domain descriptions and variable annotations, ensuring interpretability in the biological context. Our curated subset comprises 65 Boolean networks across diverse regulatory settings. For each sample, we provide 30 pairs of transitions by randomly varying the initial conditions.
(c) Structured Causal Models (SCMs). Beyond algebraic and logical dynamics, we include causal structures by deriving SCMs using functional analysis, which extracts the variable dependencies of samples from two sources: (a) our curated CDEs dataset, and (b) additional multivariate dynamical systems from the Physiome database. Each SCM is formalized as a directed graph where edges encode causal dependencies with explicit time lags. To maintain interpretability, we annotate variables with their scientific roles (e.g., ion concentrations, membrane potentials) and domains. This results in 190 SCM samples covering complex physiological and physical systems. For each sample, we sample 150 data points based on the ground truth CDE.
Stress Tests: Beyond standard in-distribution splits, SymbolBench includes evaluations under out-of-distribution (OOD) and noisy settings created by systematically varying initial conditions and control parameters outside the fitting range, and adding noise to the time series during fitting. This design mirrors real-world challenges where scientific systems rarely operate under a single controlled regime or without random external noise, thereby testing a model’s ability to generalize to unseen but realistic scenarios rather than simply interpolating observed, noise-free patterns. More details of the dataset are provided in Appendix B.
@@
Figure 2: Iterative refinement framework with hybrid design. Candidate proposals are generated using either an LLM-as-Predictor or genetic programming operations. Each round of candidates undergoes quantitative and qualitative evaluation via validation tools and an LLM-as-Judge. Scored candidates are stored in a history pool, and a context manager decides contextual information for the next round.
@@ 4.1 Proposal Generation
LLM-as-Predictor. To leverage LLMs’ embedded knowledge, prior work has explored direct equation generation, which we term LLM-as-Predictor. To evaluate both model capability and the effects of iterative refinement, context, and chain-of-thought reasoning, we design four prompting strategies for LLMs/MLLMs: (a) Naive, which generates expressions without contextual or historical input; (b) Base, which adds filtered prior expressions from the history pool; (c) Context, which further incorporates relevant context such as variable descriptions; (d) CoT, which extends Context with step-by-step reasoning, which is recorded for qualitative verification. Full prompt details are in Appendix K.
Hybrid Method. While generating symbolic structures is central to the pipeline, this step can also leverage genetic programming operations. For CDEs and BNs, this entails applying crossover on expression trees to create new candidates. As LLMs can support other components, e.g., verification, we refer to this integration as the Hybrid Method (Figure 3), as discussed in Appendix H.
Figure 3: Hybrid method.
@@ 4.2 Verification
Verification in symbolic structure discovery assesses three key aspects: numerical fitness, symbolic fitness, and context alignment. However, unlike numerical fitness, symbolic fitness and context alignment are hard to quantify straightforwardly. To address this, we introduce a dual-verification strategy combining quantitative metrics with rubric-based qualitative evaluation with LLMs-as-Judges.
@@ 4.2.1 Quantitative Verification
(a) For CDEs, we simulate time series from predicted functions and assess numerical fitness via the averaged coefficient of determination ($R^{2}$), with expression complexity measured by the number of operations. (b) For Boolean networks, we simulate transitions from predicted networks and evaluate predictive performance using the macro-averaged F1 score over $T$ transitions; expression complexity is computed as in CDEs. (c) For structural causal models, where the true generative process is unknown, we use the Conditional Independence score (CI-score) as a proxy for structural fidelity, measuring dependence between each child and its parents conditioned on the remaining parents. Lower CI-scores indicate stronger conditional independence and higher structural plausibility.
@@ 4.2.2 Qualitative Verification
To evaluate the symbolic quality of generated expressions, we apply a rubric-based scoring from 1 (poor) to 5 (excellent). Each score is assigned by an LLM judge Hao et al. (2024) following a standardized rubric with four criteria. (a) Context alignment: consistency with time series data and contextual descriptions; (b) Scientific plausibility: adherence to physical laws or domain constraints; (c) Conciseness and clarity: readability and succinctness of reasoning; (d) Logical coherence: stepwise consistency and sound derivation.
4.3 Context Managing
The Context Manager supplies relevant information to the Proposal Generator during iterative refinement. Depending on the strategy, it selects an appropriate context and maintains a history pool $\mathcal{H}$ storing prior expressions, reasoning paths, and verification scores in a structured DataFrame, with duplicates removed. We adopt a simple ranking-based approach that selects the top-$k$ highest-scoring candidates for further refinement.
@@ 5 Experiment
Table 2: Symbolic regression results for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Yellow, Orange, and Cyan mark the first, second, and third place, respectively. Statistically significant improvements over the best baseline (paired t-test, $p$ < 0.05) are underlined.